OpenStudy (cookiemonster18):

1.Suppose that the amount of algae in a pond doubles every 6 hours. If the pond initially contains 50 pounds of algae, how much algae will be in the pond after 18 hours? (1 point) A.400**** B.10,800 C.300 D.600

OpenStudy (shelbywyatt):

whats **** in a

OpenStudy (cookiemonster18):

Its suppose to be me telling what I think the answer is :3 I think #1 is A @ShelbyWyatt

OpenStudy (shelbywyatt):

like what is 400****

OpenStudy (shelbywyatt):

so lets try

OpenStudy (shelbywyatt):

first 6 hours from 50 to 100

OpenStudy (shelbywyatt):

next 6 hours 100to 200

OpenStudy (shelbywyatt):

last 6 hours 200to 400

OpenStudy (cookiemonster18):

Yes, so what do we do with these to solve them

OpenStudy (shelbywyatt):

so there will be 400 pounds of algae after 18 hours?

OpenStudy (shelbywyatt):

well 400/2 for the last 6 hours gets 200

OpenStudy (shelbywyatt):

200 for the 2nd 6 hours to give 100

OpenStudy (shelbywyatt):

100 for the first 6 hours to get 50

OpenStudy (cookiemonster18):

The answer is 900. Am I correct

OpenStudy (shelbywyatt):

how?

OpenStudy (shelbywyatt):

there are 3 6hours in 18 hours

OpenStudy (shelbywyatt):

50*2=100*2=200*2=400

OpenStudy (cookiemonster18):

So I was right the first time. The answer is (A.400)

OpenStudy (shelbywyatt):

yes

OpenStudy (shelbywyatt):

sorry for confusion though sorry

OpenStudy (cookiemonster18):

Can we solve one more than that's it :}

OpenStudy (shelbywyatt):

ok

OpenStudy (cookiemonster18):

Its fine :)

OpenStudy (cookiemonster18):

2.What is the order of the numbers from least to greatest? A=4.6x10^-4 B=2.4x10^-3 C= 3.5x10^5 D=6.3x10^-4 A.C<A<B<D B.D<A<C<B C.B<C<A<D D.A<D<B<C**** I think the answer is D.A<D<B<C

OpenStudy (cookiemonster18):

Am I correct @ShelbyWyatt

OpenStudy (shelbywyatt):

a is 460,000 i think

OpenStudy (shelbywyatt):

wait calculating each number

OpenStudy (cookiemonster18):

Oh okay sorry

OpenStudy (shelbywyatt):

a is -46,000

OpenStudy (shelbywyatt):

b is -2,400

OpenStudy (shelbywyatt):

c is 350,000

OpenStudy (shelbywyatt):

oh pellet

OpenStudy (shelbywyatt):

a is 0.00046

OpenStudy (shelbywyatt):

b is 0.0024

OpenStudy (shelbywyatt):

c is 350,000

OpenStudy (shelbywyatt):

d is 0.00063

OpenStudy (shelbywyatt):

yes you are correct

OpenStudy (cookiemonster18):

:) really? Thanks

OpenStudy (shelbywyatt):

yeah